[J, Q]
Alignment is making it tough for locks to last more than 10 minutes. Many (but not all) locklosses correlate with some optic drifting away, and taking all of the light with it. The other locklosses are the quick ones that seem to pop up out of nowhere; we haven't made any headway on these. We wanted to get to a state where we could just let the interferometer sit for some minutes, to explore the data, but got caught up with alignment and PRMI things.
We're finding that both ITMs experience some DC force when entering full PRFPMI lock. I will calculate the torque expected from radiation pressure + offset beam spot, especially for ITMX, where we choose the spot position to be uncontrolled by ASS.
I set up the QPD ASC servos to act in a common/differential way on the ETMs. The C1:ASC-XARM_[PIT/YAW] filter modules act on the common alignment, whereas the C1:ASC-YARM_[PIT/YAW] filter modules act on the differential alignment. This can soon be cleaned up with some model renaming to reduce confusion.
Using DC oplev values as a guide, we are hand tuning ITM alignment once the AO path is engaged and we see the DC drift occurring. Then, we set the QPD servo offsets and engage them.
In this manner, we were able to lock the interferometer at:
- Arm transmission 150 x single arm power
- POPDC indicated a recycling gain of ~5.5
- ASDC/POPDC indicated a contrast of 99.8%
- REFLDC indicated a visibility of 80%
We made the PRMI transition to 1f numerous times, but found that the sideband power fluctuations would get significantly worse after the transition.
We found that the gains that were previously used were too small by a factor of a few. There is a DC change visible in REFL165 before and after the transition (Also POP55, aka REFL55, is not DQ'd ). Really, it isn't certain that we've zero'd the offset in the CARM board either, so REFL55's zero crossing isn't necessarily more trustworthy that REFL165's. We can go back in the data and do some 2D histograming to see where in the error signal space the sideband power is maximized.
Jenne reports:
- The all RF transition succeeded 13/29 times.
- PRMI 1f transision succeeded 10/10 times.
|